OpenStudy (anonymous):

How do I solve this question? x + y + z = 12 6x - 2y - z = 16 3x + 4y + 2z = 28 Btw medal is awarded if u help me out :)

OpenStudy (anonymous):

use elimination let x + y + z = 12 be eqn 1 6x - 2y - z = 16 be eqn 2 3x + 4y + 2z = 28 be eqn 3 first thing you do, eliminate "z" add eqn 1 and 2-->new eqn will be eqn 4 add (eqn 1 times -2) + eqn 3--->new eqn will be eqn 5 then find either x or y using your two new equations.. then you can use substitution to the rest of your solution

OpenStudy (anonymous):

Lets use numbers plus ur explanation

OpenStudy (ranga):

Step 1) Can you add the first two equations and tell me what you are getting?

OpenStudy (anonymous):

^ yes, do that first

OpenStudy (anonymous):

You can also use matrices, and i think is simpler and easier. But above methods are correct also.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so what you get for eqn 4?

OpenStudy (anonymous):

7x - y = 28

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

now next step: multiply eqn 1 by -2

OpenStudy (anonymous):

K

OpenStudy (anonymous):

I got 4x + z = -8 right?

OpenStudy (anonymous):

multiply eqn 1 by -2 ??

OpenStudy (anonymous):

No sorry hold on

OpenStudy (anonymous):

-2x + 2y + 2z = 24

OpenStudy (anonymous):

watch your negatives...

OpenStudy (anonymous):

Oh right sorry

OpenStudy (anonymous):

so you'll get?

OpenStudy (anonymous):

U mean at the end?

OpenStudy (anonymous):

multiply eqn 1 by -2,

OpenStudy (anonymous):

Ok got that so its the same except its -

OpenStudy (anonymous):

yes,, so that will be?

OpenStudy (anonymous):

-2x - 2y - 2z = -24

OpenStudy (anonymous):

right:) so next step: add this eqn (-2x - 2y - 2z = -24) to eqn 3

OpenStudy (anonymous):

K

OpenStudy (anonymous):

Its x + 2y = 4 right?

OpenStudy (anonymous):

yes! so that will be our eqn 5 so you decide which variable (x or y) would you like to solve first?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Will it make any difference?

OpenStudy (anonymous):

ok, so again using eqn 4 and eqn 5 we will use elimination to eliminate "x"

OpenStudy (anonymous):

try multiplying eqn 5 by -7

OpenStudy (anonymous):

Wait slow down please

OpenStudy (anonymous):

no, it will not make any difference:)

OpenStudy (anonymous):

Ok I was looking for my 4th equation

OpenStudy (anonymous):

eqn 4: 7x-y=28 eqn 5: x + 2y = 4

OpenStudy (anonymous):

Yea got u

OpenStudy (anonymous):

did you multiply eqn 5 by -7 already?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sorry if i'm going too fast, i need to go offline now....

OpenStudy (anonymous):

I got -7x - 14y = -28

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

right:) then add that to eqn 4

OpenStudy (anonymous):

I did that

OpenStudy (anonymous):

Is y = 0?

OpenStudy (anonymous):

RIght!!!!!! (^_^) so next, uhm substitute y value to either eqn 4 or eqn 5 to solve for x

OpenStudy (anonymous):

Lol k thanx sooo much!! U should help me every time I got a question.

OpenStudy (anonymous):

lol XD can you manage now from here?

OpenStudy (anonymous):

Hmm...

OpenStudy (anonymous):

can you?

OpenStudy (anonymous):

Listen if u want to go just go ill figure it out and its an easy step dont worry :)

OpenStudy (anonymous):

Yes I can :D

OpenStudy (anonymous):

i can stay for 10 more minutes i think

OpenStudy (anonymous):

YAY! i like you, you're an easy-learner :) goodluck in your math, c ya round (^_^)

OpenStudy (anonymous):

Sure :) anytime lol

OpenStudy (anonymous):

Just one more question

OpenStudy (anonymous):

Hold up tho

OpenStudy (anonymous):

Im supposed to have 2 numbers or 3? U kno x, y, & z

OpenStudy (anonymous):

yeah, you need to find the values of (X, Y, Z)

OpenStudy (anonymous):

Oh k got it nvm thanx again :)

OpenStudy (anonymous):

I do have another question but since ur leaving ill ask someone else